Ideally, when a thermometer is used to measure the temperature of an object, the temperature of the object itself should not change. However, if a significant amount of heat flows from the object to the thermometer, the temperature will change. A thermometer has a mass of 34.0 g, a specific heat capacity of c = 815 J/(kg · C°), and a temperature of 12.0° C. It is immersed in 119 g of water, and the final temperature of the water and thermometer is 43.9° C. What was the temperature of the water before the insertion of the thermometer? (See Table 12.2 for appropriate constants.)
(answer in ° C)
Answer #1

Given that

mass m2=0.034 kg

mass m1=0.119 kg

temperature T2=12 C

finial temperature T3=43.9 C

specific heat capacity S2=815 J/kg.c

specific heat capacity of water S1=4184 J/kg.C

now we find the temperature of water

m1S1[T1-T3]=m2s2[T3-T1]

0.119*4184*[T1-43.9]=0.034*815[43.9-12]

497.896[T1-43.9]=883.949

T1-43.9=2.019

T1=45.919 C